OpenStudy (anonymous):

if triangle ABE~ triangle ACD, find the value of y. A. y= -1 B. y=1 C. y=2.4 D. y=6 Give me a second gota draw it:)

OpenStudy (anonymous):



OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

the triangles are similar, so AE/AD = EB/DC y/(y+4) = 6/10

OpenStudy (anonymous):

y=6 is that correct?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

y/(y+4) = 6/10 6/(6+4) = 6/10 6/10 = 6/10 ... works

OpenStudy (anonymous):

okay!